# Conformal Prediction Sets Improve Human Decision Making

### v1

@@ 1 Introduction
Prediction sets are a useful tool for augmenting the output of a discriminative model. Rather than outputting a single prediction, a model can output a set of predictions, which may contain zero, one, or several possible values. Generally, prediction sets are constructed to contain the values that are most likely to be correct according to the model using some heuristic notion of confidence. Perhaps the most straightforward way to construct prediction sets in the classification context is by including the top-$k$ values according to softmax outputs for a fixed $k$ (Grycko, 1993). Top-$k$ prediction sets are useful for providing alternatives and can even furnish statistical guarantees (Chzhen et al., 2021), however, they do not quantify uncertainty.
@@
Conformal prediction (Vovk et al., 2005) is a general-purpose method for transforming heuristic notions of uncertainty into rigorous ones through the use of calibrated prediction sets. Conformal sets usually are designed so that the set size indicates how confident the model is about a particular input, with larger sets signalling more uncertainty (Figure 1). The hallmark of conformal prediction is its risk control; conformal sets will only fail to contain the ground truth with a pre-specified error rate. Furthermore, conformal prediction is widely applicable since it can be used with any pre-trained model, does not rely on distributional assumptions, nor on infinite-data limits (Angelopoulos & Bates, 2021). It has been successfully applied in real-world settings for drug discovery (Eklund et al., 2015; Alvarsson et al., 2021), medical diagnosis (Zhan et al., 2020; Angelopoulos et al., 2022), time-series forecasting (Stankeviciute et al., 2021; Xu & Xie, 2021; Zaffran et al., 2022), and language modelling (Kumar et al., 2023; Ren et al., 2023).
@@ 2.1 Uncertainty Quantification
In order build more trustworthy machine learning models, it is paramount to reliably characterize the uncertainty in their predictions (Smith, 2013; Sullivan, 2015; Soize, 2017). In the case of classification models, while softmax scores of maximum likelihood predictions might seem to be intuitive proxies of uncertainty, it has been shown that modern neural networks do not intrinsically provide calibrated softmax scores (Guo et al., 2017; Minderer et al., 2021; Bai et al., 2021). At best, softmax scores are heuristic notions of uncertainty. However, there exists a plethora of uncertainty quantification techniques including Bayesian methods (Mackay, 1992; Neal, 2012; Graves, 2011; Blundell et al., 2015), Monte Carlo dropout (Gal & Ghahramani, 2016), and ensembles (Dietterich, 2000; Lakshminarayanan et al., 2017; Ashukha et al., 2020). Yet, many such methods can be limited in practice due to their computational overhead, incorrect distributional assumptions, or predisposition to specific architectures and training procedures.
@@ 2.2 Prediction Sets and Coverage
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[y\in\mathcal{C}(x)]\geq 1-\alpha,$ (1)
@@
Given an arbitrary classifier $f:\mathcal{X}\to[0,1]^{M}$ with softmax outputs, the most basic method of creating prediction sets is to select the top-$k$ outputs of $f(x)$ which we denote $\mathrm{top}_{f}(x,k)$,
@@ 2.3 Conformal Prediction
Conformal prediction (Vovk et al., 2005; Shafer & Vovk, 2008) is a prominent framework to construct prediction sets for arbitrary classification models [^3] While conformal prediction can applied to a broader class of discriminative models, we focus our study on classification. such that, for any pre-specified $\alpha$, the resulting sets provably satisfy the coverage guarantee in Eq. (1). In particular, it provides risk control – the ability to choose $\alpha$ – which is achieved by allowing the size of prediction sets to vary based on a heuristic notion of uncertainty built into $f$, such as softmax scores. First, we define a conformal score function $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ where larger conformal scores indicate worse agreement between an input $x\in\mathcal{X}$ and a class $y\in\mathcal{Y}$ according to the heuristic uncertainty notion. While improved alternatives exist (Romano et al., 2020; Angelopoulos et al., 2021), the canonical form of the conformal score for classification is $s(x,y)=1-f(x)_{y}$, where $f(x)_{y}$ is the softmax output for class $y$. Second, using a calibration dataset $\mathcal{D}_{\mathrm{cal}}$ of size $n$ drawn from $\mathbb{P}$, we compute the conformal threshold $\hat{q}$ as the $\tfrac{\lceil{(n+1)(1-\alpha)}\rceil}{n}$ quantile of the conformal scores $s_{i}=s(x_{i},y_{i})$ on $\mathcal{D}_{\mathrm{cal}}$. Since this quantile is computed from pairs of input data $x_{i}$ and corresponding labels $y_{i}$, it can be understood as a calibrated conformal score for the model. Finally, $\hat{q}$ is used to generate prediction sets that contain all classes $y$ such that the conformal score is below the threshold $\hat{q}$, [^4] See sketch in Figure 1 where $f(x)_{y}>1-\hat{q}$ is used in lieu of $s(x,y)<\hat{q}$.
@@
Conformal prediction sets will obey the coverage guarantee in Eq. (1) for test data $\mathcal{D}_{\mathrm{test}}$ with the sole assumption being that the calibration and test data are exchangeable (Vovk et al., 1999), a weaker assumption than being i.i.d. Hence, unlike other methods for uncertainty quantification, conformal prediction does not rely on assumptions about the underlying model $f$, its training data, nor on distributional assumptions about the data (Angelopoulos & Bates, 2021).
Importantly, conformal prediction sets are calibrated to be larger when the model has greater uncertainty in its predictions. Compared to the fixed-sized nature of top-$k$ sets, well-designed conformal sets will have a smaller average set size for the same coverage (Angelopoulos et al., 2021). This is a potentially useful feature for expediting downstream human decision making, although one that has not been scientifically verified in prior work.
@@ 2.4 Human-in-the-Loop Conformal Prediction
While conformal prediction provides statistically rigorous guarantees, its applicability in real-world scenarios is limited since it may require a post-hoc mechanism to convert prediction sets into single actionable outcomes. A small number of prior works have explored using conformal prediction with human decision makers. Straitouri et al. (2023) proposed showing conformal prediction sets to human experts who are restricted to only select an option from the set. They developed a search method to find an $\alpha$ that maximizes the expert’s accuracy. Babbar et al. (2022) proposed a method for learning a policy that is attuned to a given human expert, and defers automated model decisions when it expects the human is likely to be more accurate. By deferring examples where $|\mathcal{C}_{\hat{q}}|$ would be large and presumably less useful to a human, the average set size of non-deferred examples can be reduced for the same coverage level.
Both of these works are based on the assumption that humans can effectively make use of prediction sets to improve some aspect of their decision making. However, such assumptions have not been supported by scientific evidence. Straitouri et al. (2023) did not conduct any tests involving humans, instead creating algorithmic “experts”, and using existing datasets that were human-labelled. Babbar et al. (2022) did recruit humans in two experiments. The first compared the usefulness of top-1 predictions to conformal sets, using 30 participants who were each shown the same 18 images. These 18 images were hand-selected by the experimenters to show cases where the top-1 prediction was incorrect, but the conformal set contained the true label. Similarly, a second experiment compared conformal sets to the proposed deferral scheme, using 60 participants each shown the same set of 15 examples. Once again, the 15 examples were hand-selected by the experimenters to highlight cases where the non-deferred sets were smaller in size. Due to such data manipulations, the tests cannot be considered as a randomized controlled trials. In our work we provide scientific evidence in support of the notion that humans can effectively leverage prediction sets, and the uncertainty quantification that conformal prediction provides.
@@ 3 Method
Primarily, we aim to disentangle the effects of merely receiving several alternatives versus the uncertainty quantification provided by variable conformal set sizes. To this end, we ask humans to complete a challenging task, and either provide them with no assistance (control), a top-$k$ set, or a conformal prediction set generated with the RAPS algorithm (Angelopoulos et al., 2021) (see App. A.2). Both types of sets are derived from the same pre-trained model and come with (empirical) coverage guarantees. To isolate the uncertainty quantification aspects of conformal prediction from the level of coverage, we ensure that both types of predictions sets achieve the same empirical coverage. First, we select a $k$ for top-$k$, and evaluate the empirical risk $\hat{\alpha}$ from Eq. (3) using $\mathcal{D}_{\mathrm{cal}}$. Then, we compute a calibrated conformal threshold $\hat{q}$ using the same $\mathcal{D}_{\mathrm{cal}}$ and with the choice of risk tolerance $\alpha$ equal to the $\hat{\alpha}$ achieved by top-$k$. In both cases, new data $\mathcal{D}_{\mathrm{test}}$ drawn i.i.d. from the same distribution as $\mathcal{D}_{\mathrm{cal}}$ is shown to humans along with the appropriate set and coverage guarantee. [^5] We refrain from providing any other possible information, such as the model’s ordering of likely labels, or softmax scores.
@@
Our experiments are forced-choice tasks, where participants are shown a stimulus $x$ and must select one label $y$ out of all possible classes. Each participant is randomly assigned to one task and one treatment (control, top-$k$, or conformal), then is trained on their task, and finally completes $n$ trials. We collect data on their answers and response time.
@@
The tasks we employ have one-hot labels, so that correctness is binary. For the accuracy measurement, we model each combination of task and treatment as having an inherent difficulty level which is the human probability of correctness on each trial, $q$. Given that each person completes $n$ binary trials with success probability $q$, the accuracy distribution is binomial, and each participant of the study is represented as one random sample from $B(n,q)$. For large $n$ this distribution approximates a Gaussian with $\mu=nq$. Hence, we can perform Welch’s $t$-test (Welch, 1947) to test the null hypothesis that two treatments have the same inherent difficulty via $\mu$ [^6] For accuracy, we perform one-sided $t$-tests with the prior expectation that prediction sets are more helpful than the control, and that conformal sets are more helpful than top-$k$ sets due to their smaller average size. If one treatment is measured as having larger $\mu$, we can conclude that said treatment improves human accuracy.
@@
For measurement of decision making speed, we capture the time between presentation of a stimulus and when the participant enters their response. For every response, participants are unconstrained in the amount of time they spend deciding on their answer, but have a monetary incentive to complete the test expediently since compensation is primarily a fixed amount. [^7] This payment method is enforced by the participant recruitment platform Prolific (Prolific, 2024). We do not model a particular prior distribution of response times, instead relying on the central limit theorem. Each participant spends time on $n$ trials, and we treat their average response time as a single observation. For large sample sizes the observations are expected to be normally distributed, and so we once again apply Welch’s $t$-test to test the null hypothesis that the populations for two treatments have the same average response time.
@@ 4.1 Tasks, Datasets, and Models
Image classification, a mainstay application of computer vision, is widely performed by humans in critical settings every day. For instance, radiologists diagnose diseases by classifying X-ray images, and could potentially improve the accuracy and speed of diagnosis by leveraging machine learning classifiers (Choy et al., 2018; Agarwal et al., 2023). As a representative image classification task we used ObjectNet (Barbu et al., 2019), a dataset of common objects photographed from intentionally difficult viewpoints. We selected the 20 most common classes to reduce the number to a manageable level for humans, balanced the selected classes using stratified sampling, then split the dataset into $\mathcal{D}_{\mathrm{cal}}$ and $\mathcal{D}_{\mathrm{test}}$ maintaining class balance. We used CLIP ViT-L/14 (Radford et al., 2021) as a zero-shot classifier.
@@
Sentiment analysis is crucial in human interactions. For instance, social media users routinely need to classify sentiments in posts, comments, and reviews to understand context and communicate effectively. As an example sentiment analysis task, we used GoEmotions (Demszky et al., 2020), a dataset of Reddit comments in English with annotations of sentiment categories. GoEmotions comes with a training, validation, and test set. We selected the 10 most common classes based on the validation dataset, and balanced the selected classes using stratified sampling for the validation and test set separately. We used the remaining validation datapoints as $\mathcal{D}_{\mathrm{cal}}$ and test datapoints as $\mathcal{D}_{\mathrm{test}}$. For the model we selected a RoBERTa-Base (Liu et al., 2019) model fine-tuned on the GoEmotions training set (Lowe, 2024).
Named Entity Recognition
Humans regularly perform named entity recognition (NER) in daily life as we encounter unfamiliar proper names and use context to infer what type of entity that name represents. In the financial industry, human annotators are often employed to locate and extract entities from legal documents. Humans may rely on model recommendations to expedite extraction, but are still ultimately responsible for ensuring correctness. We created a task based on NER using the Few-NERD dataset of sentences from Wikipedia in English (Ding et al., 2021), with each word annotated as a named entity class. For our task, a single entity was selected from every sentence as the classification target. We selected the 20 most common classes in the validation dataset, and balanced the selected classes using stratified sampling for the validation and test sets separately. We used the remaining validation datapoints as $\mathcal{D}_{\mathrm{cal}}$ and test datapoints as $\mathcal{D}_{\mathrm{test}}$. For the model we used a SpanMarker RoBERTa-Large fine-tuned on the Few-NERD training set (Aarsen, 2023, 2024).
For all three tasks we used $k=3$ to compare the usefulness of top-$k$ sets and conformal sets. A summary of model performance as percentages is given in Table 1. We show the empirical risk $\hat{\alpha}$ for top-3 sets computed on $\mathcal{D}_{\mathrm{cal}}$, which is then used as $\alpha$ for conformal calibration. We also show the top-1 and top-3 accuracy, empirical conformal set coverage, and average conformal set size on $\mathcal{D}_{\mathrm{test}}$.

### v3

@@ 1 Introduction
Prediction sets are a useful tool for augmenting the output of a discriminative model. Rather than outputting a single prediction, a model can output a set of predictions, which may contain zero, one, or multiple possible values. Generally, prediction sets are constructed to contain the values that are most likely to be correct according to the model using some heuristic notion of confidence. Perhaps the most straightforward way to construct prediction sets in the classification context is by including the top-$k$ values according to softmax outputs for a fixed $k$ (Grycko, 1993). Top-$k$ prediction sets are useful for providing alternatives and can even furnish statistical guarantees (Chzhen et al., 2021); however, they do not quantify uncertainty.
@@
Conformal prediction (Vovk et al., 2005) is a general-purpose method for transforming heuristic notions of uncertainty into rigorous ones through the use of calibrated prediction sets. Conformal sets usually are designed so that the set size indicates how confident the model is about a particular input, with larger sets signalling more uncertainty (Figure 1). The hallmark of conformal prediction is its risk control; conformal sets will only fail to contain the ground truth with a pre-specified error rate. Furthermore, conformal prediction is widely applicable since it can be used with any pre-trained model, and does not rely on distributional assumptions, nor on infinite-data limits (Angelopoulos & Bates, 2021). It has been successfully applied in real-world settings for drug discovery (Eklund et al., 2015; Alvarsson et al., 2021), medical diagnosis (Zhan et al., 2020), time-series forecasting (Stankeviciute et al., 2021; Xu & Xie, 2021; Zaffran et al., 2022), and language modelling (Kumar et al., 2023; Ren et al., 2023).
@@ 2.1 Uncertainty Quantification
In order build more trustworthy machine learning models, it is paramount to reliably characterize the uncertainty in their predictions (Smith, 2013; Sullivan, 2015; Soize, 2017). In the case of classification models, while the softmax scores of maximum likelihood predictions might seem to be intuitive proxies of uncertainty, it has been shown that modern neural networks do not intrinsically provide calibrated softmax scores (Guo et al., 2017; Minderer et al., 2021; Bai et al., 2021). At best, softmax scores are heuristic notions of uncertainty. However, there exists a plethora of uncertainty quantification techniques including Bayesian methods (Mackay, 1992; Neal, 2012; Graves, 2011; Blundell et al., 2015), Monte Carlo dropout (Gal & Ghahramani, 2016), and ensembles (Dietterich, 2000; Lakshminarayanan et al., 2017; Ashukha et al., 2020). Yet, many such methods can be limited in practice due to their computational overhead, incorrect distributional assumptions, or predisposition to specific architectures and training procedures.
@@ 2.2 Prediction Sets and Coverage
$\mathbb{P}[y\in\mathcal{C}(x)]\geq 1-\alpha,$ (1)
@@
Given an arbitrary classifier $f\!:\mathcal{X}{\to}[0,1]^{M}$ with softmax outputs, the simplest method of creating prediction sets is to select the top-$k$ outputs of $f(x)$, denoted $\mathrm{top}_{f}(x,k)$,
@@ 2.3 Conformal Prediction
Conformal prediction (Vovk et al., 2005; Shafer & Vovk, 2008) is a prominent framework to construct prediction sets for arbitrary classification models [^3] While conformal prediction can be applied to a broader class of discriminative models, we focus our study on classification. such that, for any pre-specified $\alpha$, the resulting sets provably satisfy the coverage guarantee in Equation 1. In particular, it provides risk control – the ability to choose $\alpha$ – which is achieved by allowing the size of prediction sets to vary based on a heuristic notion of uncertainty built into $f$, such as softmax scores. First, we define a conformal score function $s\!:\mathcal{X}{\times}\mathcal{Y}{\to}\mathbb{R}$ where larger conformal scores indicate worse agreement between an input $x\in\mathcal{X}$ and a class $y\in\mathcal{Y}$ according to the heuristic uncertainty notion. While improved alternatives exist (Romano et al., 2020; Angelopoulos et al., 2021), the canonical form of the conformal score for classification is $s(x,y)=1{-}f(x)_{y}$, where $f(x)_{y}$ is the softmax output for class $y$. Second, using a calibration dataset $\mathcal{D}_{\mathrm{cal}}$ of size $n$ drawn from $\mathbb{P}$, we compute the conformal threshold $\hat{q}$ as the $\tfrac{\lceil{(n+1)(1-\alpha)}\rceil}{n}$ quantile of the conformal scores $s_{i}=s(x_{i},y_{i})$ on $\mathcal{D}_{\mathrm{cal}}$. Since this quantile is computed from pairs of input data $x_{i}$ and corresponding labels $y_{i}$, it calibrates the conformal score for the model. Finally, $\hat{q}$ is used to generate prediction sets that contain all classes $y$ such that the conformal score is below the threshold $\hat{q}$, [^4] See Figure 1 where $f(x)_{y}{>}1-\hat{q}$ is used in lieu of $s(x,y){<}\hat{q}$.
@@
Conformal prediction sets will obey the coverage guarantee in Equation 1 for test data $\mathcal{D}_{\mathrm{test}}$ with the sole assumption being that the calibration and test data are exchangeable (Vovk et al., 1999), a weaker assumption than being i.i.d. Hence, unlike other methods for uncertainty quantification, conformal prediction does not rely on assumptions about the underlying model $f$, its training data, nor on distributional assumptions about the data (Angelopoulos & Bates, 2021).
Importantly, conformal prediction sets are calibrated to be larger when the model has greater uncertainty in its predictions. Compared to the fixed-sized nature of top-$k$ sets, well-designed conformal sets will have a smaller average set size for the same coverage (Angelopoulos et al., 2021). This is a potentially useful feature for improving downstream human decision making, although one that has not been scientifically verified in prior work.
@@ 2.4 Human-in-the-Loop Conformal Prediction
While conformal prediction provides statistically rigorous guarantees, its applicability in real-world scenarios is limited since it requires a post-hoc mechanism to convert prediction sets into single actionable outcomes. A small number of prior works have explored using conformal prediction with human decision makers. Straitouri et al. (2023) proposed restricting experts to only select an option from a conformal prediction set, and developed a search method to find an $\alpha$ that maximizes the expert’s accuracy. Babbar et al. (2022) proposed learning a policy that is attuned to a given expert, and defers automated model decisions when it expects the human is likely to be more accurate.
These works are based on the assumption that humans can effectively make use of prediction sets compared to acting without assistance. However, such assumptions have not been supported by scientific evidence. Straitouri et al. (2023) did not conduct any tests involving humans, instead creating algorithmic “experts”. Follow-up work by Straitouri & Rodriguez (2023) did involve humans, but only compared cases where humans are or are not forced to choose their answer from a conformal set. Babbar et al. (2022) recruited humans to compare the usefulness of top-1 predictions to conformal sets and their proposed deferral scheme. However, in these tests the sets of images shown to participants were hand-selected by the experimenters to highlight cases where the top-1 prediction was incorrect, or where non-deferred sets were smaller in size. Due to such data manipulations, the test data would not be exchangeable with the calibration data, invalidating coverage guarantees. In our work we provide scientific evidence in support of the notion that humans can effectively leverage prediction sets and the uncertainty quantification that conformal prediction provides.
@@ 3 Method
Primarily, we aim to disentangle the effects of merely receiving several alternatives versus the uncertainty quantification provided by variable conformal set sizes. To this end, we ask humans to complete a challenging task, and either provide them with no assistance (control), a top-$k$ set, or a conformal prediction set generated with the RAPS algorithm (Angelopoulos et al., 2021) (see Appendix A.2). Both types of sets are derived from the same pre-trained model and come with (empirical) coverage guarantees. To isolate the uncertainty quantification aspects of conformal prediction from the level of coverage, we ensure that both types of predictions sets achieve the same empirical coverage. First, we select a $k$ for top-$k$, and evaluate the empirical risk $\hat{\alpha}$ from Equation 3 using $\mathcal{D}_{\mathrm{cal}}$. Then, we compute a calibrated conformal threshold $\hat{q}$ using the same $\mathcal{D}_{\mathrm{cal}}$ and with the choice of risk tolerance $\alpha$ equal to the $\hat{\alpha}$ achieved by top-$k$. In both cases, new data $\mathcal{D}_{\mathrm{test}}$ drawn i.i.d. from the same distribution as $\mathcal{D}_{\mathrm{cal}}$ is shown to humans along with the appropriate set and coverage guarantee. [^5] We refrain from providing any other possible information, such as the model’s ordering of likely labels, or softmax scores.
@@
Our experiments are forced-choice tasks, where participants are shown a stimulus $x$ and must select one label $y$ out of all possible classes. Each participant is randomly assigned to one task and one treatment (control, top-$k$, or conformal), then is trained on their task, and finally completes $m$ trials. We collect data on their answers and response times.
@@
The tasks we employ have one-hot labels, so that correctness is binary. Accuracy is computed as the fraction of the $m$ trials that were answered correctly. For each of $N$ participants in a treatment, we consider their accuracy over $m$ trials as one observation; for large $N$ we expect the mean accuracy follows a Gaussian distribution by the central limit theorem. Also, since each participant takes only one test, observations are independent and unpaired. Hence, we can perform Welch’s $t$-test (Welch, 1947) to test the null hypothesis that two treatments have the same mean accuracy. [^6] For accuracy, we perform one-sided $t$-tests with the prior expectation that top-$k$ sets are more helpful than the control, and that conformal sets are more helpful than top-$k$ sets due to their smaller average size.
@@
For measurement of decision making speed, we capture the time between presentation of a stimulus and when the participant enters their response. For every response, participants are unconstrained in the amount of time they spend deciding on their answer, but have a monetary incentive to complete the test expediently since compensation is primarily a fixed amount as enforced by the recruitment platform (Prolific, 2024). We treat each participant’s total response time over $m$ trials as a single observation, and again apply Welch’s $t$-test for the null hypothesis that two treatments have the same mean response time.
Finally, we provide effect sizes computed with Cohen’s $d$ for equal sample sizes: $d=(\overline{X}_{1}-\overline{X}_{2})/{\sqrt{(s_{1}^{2}+s_{2}^{2})/2}}$. Here, $\overline{X}_{i}$ is the sample mean over observations for treatment $i$, and the sample variance is $s_{i}^{2}=\frac{1}{N-1}\sum_{j}^{N}(X^{j}_{i}-\overline{X}_{i})^{2}$, using observations $X^{j}_{i}$ from treatment $i$ (Cohen, 2013).
@@ 4.1 Tasks, Datasets, and Models
Image classification, a mainstay application of computer vision, is widely performed by humans in critical settings every day. For instance, radiologists diagnose diseases by classifying X-ray images, and could potentially improve the accuracy and speed of diagnosis by leveraging machine learning (Choy et al., 2018; Agarwal et al., 2023). As a representative image classification task we used ObjectNet (Barbu et al., 2019), a dataset of common objects photographed from intentionally difficult viewpoints. We selected the 20 most common classes to reduce the number to a manageable level for humans, balanced the selected classes using stratified sampling, then split the dataset into $\mathcal{D}_{\mathrm{cal}}$ and $\mathcal{D}_{\mathrm{test}}$ maintaining class balance. We used CLIP ViT-L/14 (Radford et al., 2021) as a zero-shot classifier.
@@
Sentiment analysis is crucial in human interactions. For instance, social media users routinely classify sentiments in posts, comments, and reviews to understand context and communicate effectively. As a prototypical task, we used GoEmotions (Demszky et al., 2020), a dataset of Reddit comments in English with annotations of sentiment categories. We selected the 10 most common classes based on the validation dataset, and balanced the classes using stratified sampling for the validation and test set separately to form $\mathcal{D}_{\mathrm{cal}}$ and $\mathcal{D}_{\mathrm{test}}$. For the model we selected a RoBERTa-Base (Liu et al., 2019) fine-tuned on the GoEmotions training set (Lowe, 2024).
Named Entity Recognition (NER)
Humans regularly perform NER in daily life as we encounter unfamiliar proper names and use context to infer what type of entity that name represents. In the financial industry, human annotators are often employed to locate and extract entities from legal documents, but may rely on model recommendations to expedite extraction. We used the Few-NERD dataset of sentences from Wikipedia in English (Ding et al., 2021), with each word annotated as a named entity class. For our task, a single entity was selected from every sentence as the classification target. We selected the 10 most common classes based on the validation dataset, and balanced the classes using stratified sampling for the validation and test set separately to form $\mathcal{D}_{\mathrm{cal}}$ and $\mathcal{D}_{\mathrm{test}}$. For the model we used a SpanMarker RoBERTa-Large fine-tuned on the Few-NERD training set (Aarsen, 2023, 2024).
For all three tasks we used $k=3$ to compare the usefulness of top-$k$ sets and conformal sets. A summary of model performance as percentages is given in Table 1. We show the empirical risk $\hat{\alpha}$ for top-$k$ sets computed on $\mathcal{D}_{\mathrm{cal}}$, which is then used as $\alpha$ for conformal calibration. We also show for $\mathcal{D}_{\mathrm{test}}$ the top-1 and top-3 accuracy of the models, empirical conformal set coverage, and average conformal set size.